OpenStudy (anonymous):

Let f(x) = 3x - 10. Find the value of f(-3). Answer 21 -1 -19

OpenStudy (whpalmer4):

\[f(x) = 3x-10\]Find \(f(-3)\) To do so, simply substitute \((-3)\) wherever you see \(x\), and do the arithmetic. For example, \[f(5) = 3(5)-10 = 15-10 = 5\]